OpenStudy (anonymous):

Select the ordered pair from the choices below that satisfies the following system of equations: 3x − 4y = 18 5y = x − 28

OpenStudy (anonymous):

Whoever finds the asnser first gets a medal!

OpenStudy (mathmate):

You have not given the choices. It is easier with the choices because we don't need to solve the equations.

OpenStudy (anonymous):

divide second equation by 5 \[\Large y= \frac{x-28}{5}\]

OpenStudy (anonymous):

put the above value in the first equation for y \[\Large 3x-4(\frac{x-28}{5})=18\] simplify this and solve for x

OpenStudy (anonymous):

@CrisYB can you tell me value of x now ?

OpenStudy (anonymous):

if all you need is the answer, no need to post here an give medals http://www.wolframalpha.com/input/?i=3x+%E2%88%92+4y+%3D+18%2C+5y+%3D+x+%E2%88%92+28

OpenStudy (anonymous):

if you want help figuring out how to do it yourself, this would be a good place to ask, but if you only need the answer just type it in